wiki:UC20170109

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Phonecon, 1/09/17

Attending: Nancy Chanover, Mark Klaene, Rene Walterbos, Jon Holtzman, Russet McMillan, John Bally, Mary Beth Kaiser, Rachel Kuzio de Naray, Anne Verbiscer, Scott Anderson, Chip Kolbulnicky, Jason Lozo, Ben Williams, Bill Ketzeback, John Tobin, and Bruce Gillespie


User feedback and comments from institutional representatives:

  • Washington - Scott had nothing new to report.
  • Georgia State - Rachel said she appreciated getting some of the open nights.
  • Seattle - No report.
  • NMSU - Jon had nothing new to report.
  • JHU - Mary Beth had nothing new to report.
  • Oklahoma - John T. had nothing new to report.
  • Colorado - John B. had nothing new to report.
  • Wyoming - Chip had nothing new to report.
  • Virginia - Anne had nothing new to report.
  • NAPG - No report.
  • Adler - No report.

Telescope and Instruments Report:

Mark noted that there was a recent software problem with the secondary mirror actuation, currently being worked on. ARCTIC is in the lab disassembled, getting its cold straps modified. Bill said he expects it to be back in service later this week. The operating temperature of the camera has been changed from 155K to 185K. Mark’s detailed report follows below.

3.5-m Telescope and Instruments Highlights, 12/7/16 through 1/4/17

0) Overview

Winter has come but a La Niña year is expected so less than average snow is expected. However, high humidity, threatening precipitation, and clouds have plagued us so far this winter. Telescope has been performing well. We made it through the holiday break with minimal problems but also minimal on-sky time due to poor weather.

1) Telescope - Telescope was operational with no issues.

2) Instruments

DIS blue and red cameras have been operational. The red camera has showed some indication of returning contamination but still at a minimal amount. A Conceptual PDR for several possible upgrades is being planned for the next month or so.

TripleSpec has produced only corrupted images during this last period. We are planning for a temperature cycle early 2017, and if that doesn’t work then a UVa servicing mission will occur in the summer. For all practical matters TSpec is not available.

Agile has been operational this period.

Echelle has been operational with no problems.

NIC-FPS is operational.

ARCTIC is off line for improvements to its thermal management. The dewar is running too cold for the Leach pre-amp electronics. This causes the loss of a quadrant when restarted. The plan is to remove several layers of thermal strapping to allow the detector to run warmer with the heater on during the cold winter months. We are also going to work on improving the diffuser rotator bearing during this period.

3) ARCSAT - Has been performing nominally.


Q1 3.5m schedule updates:

Nancy said that much of the open and DD time has been allocated, but there is still a small number of slots available — e.g., there are 3 hours of B-half time still open for 28 March. If interested, contact Nancy.


Q1 ARCSAT schedule update:

Nancy noted that the Q1 ARCSAT schedule was published. Russet added that 10 April will be the start date for the next scheduling quarter.


Update on new DIS slits:

Bill said that the new 5 arcsec slit that was made by Jason Lozo has been well-received by the first users. Jason has also procured a 20 arcsec slit that is now available for use at the telescope. We have other new slits in queue for fabrication in three sizes, 0.9, 1.5, and 2.0 arcsec widths. Jason said that if there are other slit widths or configuration that users need, speak up now while we have the production pipeline is in place. Nancy added that any special slit fabrication requests (beyond the aforementioned complement: 0.9, 1.5, 2.0, 5.0, and 20 arcsec) should be forwarded to her.

ACTION:

  • User committee members to poll their users to see if other new slit configurations are needed at the telescope, and get back to Nancy by 19 January.

Closure on issue of DIS B1000 grating:

Nancy and Mark confirmed last month’s decision regarding the default blue grating for DIS. We will not change from the default B1200 grating until the upcoming DIS review is held, and new policies are established regarding grating changes.


Update on DIS upgrades PDR:

Nancy announced that the DIS upgrades PDR will be held on 25 January at UWashington. The review panel is being formed, and the review documents are nearly ready.


WIRO time trades redux:

Nancy asked if there was feedback from last month’s action item on this topic. Anne asked about the WIRO seeing statistics, and Chip said that the site is 1.5-1.8 arcsec fwhm in the visible. Scott said he was waiting for replies, and Mary Beth noted that it was difficult to get any user feedback over the holidays. Nancy decided to keep the action open for another month.


Continued discussion of ARCTIC filters:

Bill said that the G and R Sloan filters will be ordered for ARCTIC. Bill noted that a 30 Å H-alpha filter would be preferable because of the location of nearby nitrogen lines. Because we would desire a narrower bandpass we would not benefit from a price break at this time on another order that the company is filling for a European customer. Also, it was noted that in order to take full advantage of the larger filters, we will need to modify the NA2 stray-light baffle system; in the meantime, the existing 3-inch filters work adequately. It was also mentioned that there are 4-inch round H-alpha and H-alpha-redshift filters available from NMSU that are currently at APO. Getting a new SiII filter will be a lower priority, since we have a 3-inch square one already. Therefore any further decision on full format narrow band filters will be tabled until the NA2 baffle project is completed.


Miscellaneous:

Russet has made minor updates to the schedule forms that consolidate information sections, change the headings to be more obvious, and have new descriptive text. Nancy and Ben will review the updated schedule forms before they are launched.


ACTION ITEMS:

open from previous meetings:

  • Committee members to query their users on their interest in using the WIRO facility.

Status: Hold open for another month, see above.

new from this meeting:

  • User committee members to poll their users to see if other new slit configurations are needed at the telescope, and get back to Nancy by 19 January.

Next meeting:

The next Users Committee phonecon will be on Monday 6 February at 1030 MST — the agenda and other materials will be sent to committee members during the preceding week.

Note: See TracWiki for help on using the wiki.